The blood level of glucose after eating a hamburger may rise to 130mg/100mL.
(i). Express this concentration in terms of molarity.
(ii). If the blood volume of a male subject is normal at 85mL/kg, calculate, for a total body mass of 80kg, the total mass of glucose in the blood.
1
Expert's answer
2020-08-02T15:51:55-0400

Molarity=mole of solute/volume of solution in L

Mole of glucose=130mg*1g/1000mg*1mole/180.16g=7.22*10^-4 mol

Volume of solution in L=100mL*1L/1000mL=0.1L

Molarity=7.22*10^-4mol/0.1L=7.22*10^-3mol/L


85mL/kg*80kg=6800mL=6.8L

Mole of glucose=6.8L*7.22*10^-3mol/L=0.0491mol

Mass of glucose=0.0491mol*180.16g/mol=8.85g